ali painted the outside of the cabinet measure 3m x 2m x 2.5m. how much surface area did he cover if painted all except the bottom of the cabinet
1
Expert's answer
2020-09-18T16:56:43-0400

The cabinet without bottom has 5 faces: top is rectangles with sides 3m and 2m (its area is "3m\cdot 2m=6m^2"), 2 lateral fases are rectangles with sides 3m and 2.5m (their area is "2\cdot 3m\cdot 2.5m=15m^2") and 2 lateral fases are rectangles with sides 2m and 2.5m (their area is "2\cdot 2m\cdot 2.5m=10m^2").

So covered area is "6m^2+15m^2+10m^2=31m^2"